OpenStudy (anonymous):

More Indefinite Integrals ! Indefinite Integral 3 sq root (x^2) (dx) I have no idea what 3sq root (x^2) is? Help is appreciated!

OpenStudy (anonymous):

First help me rewrite the integral... \[\sqrt[3]{x^2}\] (dx)

zepdrix (zepdrix):

\[\Large\rm \int\limits \sqrt[3]{x^2}dx\]Do you mean like this? cube root? Ah ok yes.

OpenStudy (anonymous):

yes !

zepdrix (zepdrix):

We want to rewrite our root as a rational exponent. Example:\[\Large\rm \sqrt[5]{x^3}=x^{3/5}\]The `numerator` represents the degree of the `power` on x. The `denominator` represents the degree of the `root`.

OpenStudy (anonymous):

so that means...\[\sqrt[3]{x^2} = x^{2/3}\]

zepdrix (zepdrix):

Mmm good good good.

zepdrix (zepdrix):

And you apply your power rule to integrate it.

OpenStudy (anonymous):

ok! great, thank you @zepdrix :)

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

my final answer is \[\frac{ 3 }{ 5 } x ^{5/3}\]

zepdrix (zepdrix):

So when you divided by the new exponent, it flipped? Yah looks good!

OpenStudy (anonymous):

yes ! this is the new rule I learned ? Its because I want to simplify the integral

OpenStudy (anonymous):

@zepdrix can you help me with this one on how they got -5? \[\frac{ 1 }{ x^5 } (dx)\]

zepdrix (zepdrix):

Jaza Jaza Jaza -_- Oh boy, you really need to brush up on your exponent rules!

OpenStudy (anonymous):

Yes I do! >.< that's why I'm here!

zepdrix (zepdrix):

\[\Large\rm x^{-a}=\frac{1}{x^a}\]The negative tells us to take the reciprocal. Examples:\[\Large\rm \frac{1}{x^3}=x^{-3}\] \[\Large\rm \frac{1}{x^{-2}}=x^2\]

zepdrix (zepdrix):

You're only doing this to the x though. So here is another important example:\[\Large\rm \frac{2}{x^5}=2x^{-5}\]So when I say reciprocal, I mean the 1/x^5 portion. Notice that the 2 still stays on top.

OpenStudy (anonymous):

I see!! That's simply put! Thank you :)

OpenStudy (anonymous):

so that means when I integrate this I will get \[x ^{-5+1} = x ^{-4}\] and then...\[\frac{ x^{-4} }{ 4 } \]

OpenStudy (anonymous):

is this right?

zepdrix (zepdrix):

\[\Large\rm \int\limits x^{-5}dx=\frac{x^{-4}}{-4}\]Woops, missing a minus sign there.

OpenStudy (anonymous):

Ok, so now I know the minus sign doesn't "go away" if need be in the denominator

zepdrix (zepdrix):

Yah, if you wanted to, you could rewrite it in the form it was given to you.\[\Large\rm \frac{x^{-4}}{-4}=\frac{1}{-4x^4}\]To remove the negative it just switches places. If it's in the numerator, it goes to the denominator (as in this example). If it was in the denominator, it would move to the numerator position.

zepdrix (zepdrix):

Or were you talking about the other minus sign?

OpenStudy (anonymous):

the minus sign, but in the previous response you answered my next question! :)

zepdrix (zepdrix):

It doesn't really matter `where` you put the negative sign. It's just a negative one, you can float it around anywhere.\[\Large\rm \frac{-1}{2}=\frac{1}{-2}=-\frac{1}{2}\]If that's what your question is about. So I would write the answer like this probably:

zepdrix (zepdrix):

\[\Large\rm =-\frac{1}{4x^4}\]

OpenStudy (anonymous):

Oh okay!! Got it, wow thank you @zepdrix